In the GDP function Y = C + S + I, where C=consumption (a+bY), S = savings and I = investment, how are C+S+I impact when... 1. Y (real income) increases? 2. r (real interest rate) increases)?
Answer #1

GDP Function:

Y = C+S+I \Rightarrow Y = (a+bY)+S+I \Rightarrow Y(1-b) = a+S+I

(1-b)

1.

When real income Y increases,

C = a+bY increases as well

Thus C+S+I increases with an increase in Y

2.

Investment is negatively related with r.

Thus when r increases, I falls down.

Thus C+S+I decreases with an increase in r.